# Freeness of arrangements of lines and one conic with ordinary quasi-homogeneous singularities

### v4

@@ 4 Graphical realizations of some arrangements
Figure 1: Arrangement of type $(1,9;6,4,6)$
@@
Figure 2: Arrangement of type $(1,8;8,2,5)$
@@
Figure 3: Arrangement of type $(1,7;8,1,4)$
@@
Figure 4: Arrangement of type $(1,7;5,4,3)$
@@
Figure 5: Arrangement of type $(1,6;3,0,4)$
@@ 5 Combinatorial constraints on plane curves with ordinary quasi-homogeneous singularities
In this short section we want to present our proof of Theorem 1.4.
@@
We present an outline of our proof, since it is similar to [12, Theorem B]. Since ${\rm deg}\,C\geqslant 6$, then we can work with the pair $(\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C)$ that is effective, and thus we can use an orbifold Bogomolov-Miyaoka-Yau inequality in the sense of [9], namely
$(\star)\,:\quad\sum_{p\in{\rm Sing}(C)}3\bigg{(}\frac{1}{2}(\mu_{p}-1)+1-e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}\bigg{)}\leqslant\frac{5}{4}d^{2}-\frac{3}{2}d,$
where $e_{orb}(p;\mathbb{P}^{2}_{\mathbb{C}},\alpha C)$ is the local orbifold Euler number of a given singularity $p$ and $\mu_{p}$ denotes the local Milnor number of $p$. Let us recall these local orbifold Euler numbers for our selection of singularities. Using [9, Theorem 8.7, Theorem 9.4.2] and a well-known formula for local Milnor numbers of ordinary singularities, we have the following:
  • If $p\in{\rm Sing}(C)$ is a node, then $\mu_{p}=1$, and $e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}=\frac{1}{4}$.
  • If $q\in{\rm Sing}(C)$ is an ordinary triple point, then $\mu_{q}=4$, and $e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}=\frac{1}{16}$.
  • If $r\in{\rm Sing}(C)$ is an ordinary quadruple point, then $\mu_{r}=9$, and $e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}=0$ which follows from the fact that the log canonical threshold for ordinary quadruple points is equal to $\frac{1}{2}$, see Remark 2.6.
@@
$3n_{2}\cdot(1-1/4)+3n_{3}\cdot(3/2+1-1/16)+3n_{4}\cdot 4=\frac{9}{4}n_{2}+\frac{117}{16}n_{3}+15n_{4},$
@@
$\frac{9}{4}n_{2}+\frac{117}{16}n_{3}+15n_{4}\leqslant\frac{5}{4}d^{2}-\frac{3}{2}d,$
@@
I declare that there is no conflict of interest regarding the publication of this paper.
@@ Acknowledgement
I want to thank Tomasz Pełka for discussions about conic-line arrangements.

### v5

@@ 4 Graphical realizations of some arrangements
Figure 1: Arrangement of type $(9;6,4,6)$.
@@
Figure 2: Arrangement of type $(8;8,2,5)$.
@@
Figure 3: Arrangement of type $(7;8,1,4)$.
@@
Figure 4: Arrangement of type $(7;5,4,3)$.
@@
Figure 5: Arrangement of type $(6;3,0,4)$.
@@ 5 Combinatorial constraints on plane curves with ordinary quasi-homogeneous singularities
In this section we present our proof of Theorem 1.4.
@@
We present a detailed outline of our proof since it is similar to [15, Theorem B]. By the assumption $m={\rm deg}\,C\geqslant 6$, so we can work with the pair $(\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C)$ that is effective, and thus we can use an orbifold Bogomolov-Miyaoka-Yau inequality [10, Section 11], namely
$(\star)\,:\quad\sum_{p\in{\rm Sing}(C)}3\bigg{(}\frac{1}{2}(\mu_{p}-1)+1-e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}\bigg{)}\leqslant\frac{5}{4}m^{2}-\frac{3}{2}m,$
where $e_{orb}(p;\mathbb{P}^{2}_{\mathbb{C}},\alpha C)$ is the local orbifold Euler number of a given singularity $p\in{\rm Sing}(C)$ [10, Definition 3.1], and $\mu_{p}$ denotes the Milnor number of $p$. Let us recall these local orbifold Euler numbers for our selection of singularities. Using [10, Theorem 8.7, Theorem 9.4.2], we can find the following:
  • If $p\in{\rm Sing}(C)$ is a node, then $e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}=\frac{1}{4}$.
  • If $q\in{\rm Sing}(C)$ is an ordinary triple point, then $e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}=\frac{1}{16}$.
  • If $r\in{\rm Sing}(C)$ is an ordinary quadruple point, then $e_{orb}\bigg{(}p;\mathbb{P}^{2}_{\mathbb{C}},\frac{1}{2}C\bigg{)}=0$ which follows from the fact that the log canonical threshold for ordinary quadruple points is equal to $\frac{1}{2}$, see Remark 2.7.
@@
$3n_{2}\cdot(1-1/4)+3n_{3}\cdot(3/2+1-1/16)+3n_{4}\cdot 5=\frac{9}{4}n_{2}+\frac{117}{16}n_{3}+15n_{4},$
@@
$\frac{9}{4}n_{2}+\frac{117}{16}n_{3}+15n_{4}\leqslant\frac{5}{4}m^{2}-\frac{3}{2}m,$
@@
We declare that there is no conflict of interest regarding the publication of this paper.
Data Availability Statement
We do not analyse or generate any datasets, because this work proceeds within a theoretical and mathematical approach.
@@ Acknowledgement
I want to thank Tomasz Pełka for discussions about conic-line arrangements. I would like to thank the anonymous referees for all their comments and suggestions that improved the presentation of this paper.